The probability that a circuit board coming off an assembly line needs rework is 0.1. Suppose that 10
boards are tested and all boards are independent of each other, and let ? be the number of circuit boards
that need rework.
(a) Find ? ?!
(b) What is the probability that exactly 8 boards need rework?
Answer #1

X ~ Bin ( n , p)

Where n = 10 , p = 0.1

a)

E(X) = n p = 10 * 0.1 = 1

b)

Binomial probability distribution is

P(X) = nCx * px * ( 1 - p)n-x

So,

P(X = 8) = 10C8 * 0.18 * ( 1 - 0.1)2

= 0.0000